# On the effects of protection zone and directed population flux in prey-predator dynamics

### v1

@@ 1 Introduction
From the perspective of reaction-diffusion models, a pioneering contribution was made by Du and Shi [3], who introduced a prey-predator model in which only the prey are allowed to enter a protection zone, while both species undergo random diffusion. Among other things, they demonstrated, through an analysis of the global structure of coexistence steady states, that enlarging the protection zone increases the likelihood of prey survival. Since their groundbreaking work, extensive studies have been conducted on Lotka-Volterra-type systems with protection zones, providing valuable insights into the structure and bifurcation behavior of steady states. In particular, various functional responses, diffusion, and cross-diffusion mechanisms have been incorporated; see, for instance, [5, 6, 21, 22, 23, 13, 12, 14, 11] and references therein. More recently, diffusion-advection effects have been explored by Ma and Tang [19] and Tang and Chen [29], revealing new qualitative features absent in purely diffusive models. Protection zones have also been studied in connection with long-time dynamics, free-boundary problems, and additional ecological mechanisms, such as strong Allee effects and fear effects; see, for instance, [7, 28, 32, 30]. However, much of the existing literature has focused primarily on stationary problems and arguments based on comparison theorems. Consequently, relatively few studies have addressed the existence and behavior of time-dependent solutions in cases involving nonlinear diffusion where comparison theorems are inapplicable, despite their critical role in ecological dynamics.
@@ 2 Local-in-time solvability
In order to construct a local-in-time solution of (2.5), we check the applicability conditions of the existence theorem by Yagi [33, Theorem 5.5]. As structural assumptions for $A$, we list the conditions given in [33, p. 202]:
@@
We now verify that these conditions are satisfied for our problem. By a standard argument, one can check that the operator $A(\boldsymbol{U})$ is sectorial and satisfies (A1) and (A2); see, for instance, the discussion in [33, p. 244]. Moreover, (A3) is evidently fulfilled by taking $\nu=1$. To verify (A4) with $\nu=1$, note that
@@
for all $\boldsymbol{U}_{1}$, $\boldsymbol{U}_{2}\in K_{R}$ and $\boldsymbol{\varPhi}\in\mathbb{H}^{2}$. As in [33, Section 8.3], one can verify that the operator $A$, defined by (2.3), satisfies (2.6) for $\mathcal{Y}=\mathbb{H}^{1+\varepsilon_{0}}$ with $0<\varepsilon_{0}<\varepsilon_{1}$. Taking into account the correspondence between the domain of the fractional power of $A(\boldsymbol{U})$ and the interpolation spaces,
@@
valid for all $\boldsymbol{U}\in\mathbb{H}^{1+\varepsilon_{1}}$ (see [33, Section 8.2]), we deduce that (A5) holds with $\beta_{0}=(1+\varepsilon_{0})/2$ and $\beta_{1}=(1+\varepsilon_{1})/2$. Consequently, (A6) is also satisfied.
Next, we state the conditions for the nonlinear mapping $F$, as extracted from [33, p. 219]:
@@
Consequently, we may apply [33, Theorems 5.5 and 5.6] to deduce the unique existence of local-in-time solutions to (2.5).
@@
It should be noted that the corresponding solution remains nonnegative for all $t\in[0,T]$. Moreover, if the initial data are nonnegative and not identically zero, then the solution becomes strictly positive. This property follows from the parabolic maximum principle and is also discussed in [33, Section 8.4].
@@ 3 Main results on the stationary problem
where $W^{2,p}_{n}(O)=\{\,w\in W^{2,p}(O):\partial_{n}w=0\ \text{on }\partial O\,\}$ $(O=\Omega\ \mbox{or}\ \Omega_{1})$. We call $(u,v)\in X$ a solution of (3.1) (or equivalently (3.2)) if it satisfies these equations. It should be noted that, due to the discontinuity of $b\,\mathrm{1}\hskip-2.5pt\mathrm{l}_{\Omega_{1}}$ across the interface $\partial\Omega_{0}$, such a solution $(u,v)\in X$ satisfies $v\in C^{2}(\overline{\Omega}_{1})$, whereas $u$ does not necessarily belong to $C^{2}(\overline{\Omega})$. A pair $(u,v)\in X$ is called a positive solution of (3.1) (equivalently (3.2)) if $u>0$ in $\Omega$ and $v>0$ in $\Omega_{1}$. In spite of the discontinuity of $b\,\mathrm{1}\hskip-2.5pt\mathrm{l}_{\Omega_{1}}$, it follows from the Sobolev embedding theorem and the Bony strong maximum principle (e.g., [1], [16, Theorem 7.1]) that any nonnegative solution $(u,v)\in X$ whose components are both nontrivial must actually be strictly positive. Such a positive solution corresponds to a biological coexistence of prey and predator within the refuge-type domain $\Omega_{1}$. The main results involve the global bifurcation structure of positive solutions of (3.1) (equivalently (3.2)), together with their asymptotic behavior as the nonlinear diffusion parameter $\alpha\to\infty$.
@@
The curve $\lambda=\sigma_{1}(b\mu\mathrm{1}\hskip-2.5pt\mathrm{l}_{\Omega_{1}},\Omega)$ for $\mu\geq 0$ represents the set of parameter values $(\lambda,\mu)$ where positive solutions bifurcate from the semitrivial solution $(0,\mu)$; this fact will be established in Theorem 3.2 below. The following lemma, derived from [3, Theorem 2.1] (see also [21, Lemma 2.1]), shows that the curve $\lambda=\sigma_{1}(b\mu\mathrm{1}\hskip-2.5pt\mathrm{l}_{\Omega_{1}},\Omega)$ is monotone increasing and uniformly bounded for $\mu\geq 0$, see Figure 2.
@@ 4.1 A priori estimates and the nonexistence region for positive stationary solutions
Let $x_{0}\in\overline{\Omega}_{1}$ be a maximum point of $v$, so that $v(x_{0})=\max_{\overline{\Omega}_{1}}v$. Then applying the maximum principle (e.g., [17, Proposition 2.1]) to the second equation of (3.2), we have
@@
In addition, we note that $\sigma_{1}(\xi\mathrm{1}\hskip-2.5pt\mathrm{l}_{\Omega_{1}})$ is continuously differentiable for $\xi>0$. The proof of the $C^{1}$-regularity can be found, for instance, in Yamada [34, Proposition 1.1 (iii)]. In his work, a continuous function, depending continuously on the parameter, is introduced as a potential function. Nevertheless, the proof remains valid in the above setting, where the potential is given by the step function $\xi\mathrm{1}\hskip-2.5pt\mathrm{l}_{\Omega_{1}}$. It follows from the chain rule that $K(\lambda,\mu,\alpha)$ is continuously differentiable. Since
@@
We remark that in the case of $\alpha=0$, (3.2) has no positive solution if $\lambda\leq\sigma_{1}(b\mu\mathrm{1}\hskip-2.5pt\mathrm{l}_{\Omega_{1}})$ (see [3, Theorem 2.1] and [21, Lemmas 2.1 and 3.3]).
@@ 4.2 Bifurcation structure of positive stationary solutions
and thus $v_{\infty}=\mu$ in $\overline{\Omega}_{1}$. Then $(\lambda_{\infty},u_{\infty},v_{\infty})=(\sigma_{1}(b\mu\mathrm{1}\hskip-2.5pt\mathrm{l}_{\Omega_{1}}),0,\mu)$ must hold by Lemma 4.5. But this contradicts (4.19) and (4.21). Hence the assertion (4.20) holds true. Thanks to the global bifurcation theorem [15, 25, 26], $\varGamma_{1}$ satisfies one of the followings:
@@
may exhibit a discontinuous jump of size $bv_{k}(x)$ across the interface $\partial\Omega_{0}\,(\subset\Omega)$. Nevertheless, even in this situation, one can invoke the generalized Harnack inequality due to Lou and Ni [18, Lemma 3.1], whose proof relies on the classical framework of Stampacchia [27], in particular on the Harnack inequality established there for divergence-form operators with merely bounded measurable coefficients and potential functions in $L^{N/2}(\Omega)$.

### v2

@@ 1 Introduction
From the viewpoint of reaction–diffusion models with protection zones (or refuges), it is important to note that the incorporation of protected regions into spatial population dynamics predates the later prey–predator literature and goes back at least to the pioneering works of López-Gómez [21] and López-Gómez and Sabina de Lis [19]. In these early contributions, spatially heterogeneous structures with protected patches were introduced within reaction–diffusion frameworks in a systematic way, and their effects on persistence/coexistence and long–term dynamics were already analyzed. This foundational line of research was subsequently developed, unified, and placed in a broader conceptual setting in the monograph of López-Gómez [23], which revisits the evolution of the protection–zone theory for parabolic equations and systems. Later, Du and Shi [3] proposed a specific prey–predator model in which only the prey are allowed to enter a protection zone, while both species undergo random diffusion. Within this particular setting, they analyzed the global structure of coexistence steady states and, in line with the earlier insights developed in [21, 19, 23], showed that enlarging the protection zone can increase the likelihood of prey survival.
Since then, extensive studies have been conducted on Lotka–Volterra-type systems with protection zones/refuges, providing valuable insights into the structure and bifurcation behavior of steady states. In particular, various functional responses, diffusion, and cross-diffusion mechanisms have been incorporated; see, for instance, [5, 6, 29, 30, 31, 14, 13, 15, 12] and references therein. More recently, diffusion–advection effects have been explored by Ma and Tang [27] and Tang and Chen [37], revealing new qualitative features absent in purely diffusive models. Protection zones have also been studied in connection with long-time dynamics, free-boundary problems, and additional ecological mechanisms, such as strong Allee effects and fear effects; see, for instance, [7, 36, 40, 38]. We also refer to the monograph [23] for a systematic account of metasolutions and the role of protected zones in the dynamics of spatially heterogeneous parabolic equations and systems. However, much of the existing literature has focused primarily on stationary problems and arguments based on comparison theorems. Consequently, relatively few studies have addressed the existence and behavior of time-dependent solutions in cases involving nonlinear diffusion where comparison theorems are inapplicable, despite their critical role in ecological dynamics.
@@ 2 Local-in-time solvability
In order to construct a local-in-time solution of (2.5), we check the applicability conditions of the existence theorem by Yagi [41, Theorem 5.5]. As structural assumptions for $A$, we list the conditions given in [41, p. 202]:
@@
We now verify that these conditions are satisfied for our problem. By a standard argument, one can check that the operator $A(\boldsymbol{U})$ is sectorial and satisfies (A1) and (A2); see, for instance, the discussion in [41, p. 244]. Moreover, (A3) is evidently fulfilled by taking $\nu=1$. To verify (A4) with $\nu=1$, note that
@@
for all $\boldsymbol{U}_{1}$, $\boldsymbol{U}_{2}\in K_{R}$ and $\boldsymbol{\varPhi}\in\mathbb{H}^{2}$. As in [41, Section 8.3], one can verify that the operator $A$, defined by (2.3), satisfies (2.6) for $\mathcal{Y}=\mathbb{H}^{1+\varepsilon_{0}}$ with $0<\varepsilon_{0}<\varepsilon_{1}$. Taking into account the correspondence between the domain of the fractional power of $A(\boldsymbol{U})$ and the interpolation spaces,
@@
valid for all $\boldsymbol{U}\in\mathbb{H}^{1+\varepsilon_{1}}$ (see [41, Section 8.2]), we deduce that (A5) holds with $\beta_{0}=(1+\varepsilon_{0})/2$ and $\beta_{1}=(1+\varepsilon_{1})/2$. Consequently, (A6) is also satisfied.
Next, we state the conditions for the nonlinear mapping $F$, as extracted from [41, p. 219]:
@@
Consequently, we may apply [41, Theorems 5.5 and 5.6] to deduce the unique existence of local-in-time solutions to (2.5).
@@
It should be noted that the corresponding solution remains nonnegative for all $t\in[0,T]$. Moreover, if the initial data are nonnegative and not identically zero, then the solution becomes strictly positive. This property follows from the parabolic maximum principle and is also discussed in [41, Section 8.4].
@@ 3 Main results on the stationary problem
where $W^{2,p}_{n}(O)=\{\,w\in W^{2,p}(O):\partial_{n}w=0\ \text{on }\partial O\,\}$ $(O=\Omega\ \mbox{or}\ \Omega_{1})$. We call $(u,v)\in X$ a solution of (3.1) (or equivalently (3.2)) if it satisfies these equations. It should be noted that, due to the discontinuity of $b\,\mathrm{1}\hskip-2.5pt\mathrm{l}_{\Omega_{1}}$ across the interface $\partial\Omega_{0}$, such a solution $(u,v)\in X$ satisfies $v\in C^{2}(\overline{\Omega}_{1})$, whereas $u$ does not necessarily belong to $C^{2}(\overline{\Omega})$. A pair $(u,v)\in X$ is called a positive solution of (3.1) (equivalently (3.2)) if $u>0$ in $\Omega$ and $v>0$ in $\Omega_{1}$. In spite of the discontinuity of $b\,\mathrm{1}\hskip-2.5pt\mathrm{l}_{\Omega_{1}}$, it follows from the Sobolev embedding theorem and the Bony strong maximum principle (e.g., [1], [22, Theorem 7.1]) that any nonnegative solution $(u,v)\in X$ whose components are both nontrivial must actually be strictly positive. Such a positive solution corresponds to a biological coexistence of prey and predator within the refuge-type domain $\Omega_{1}$. The main results involve the global bifurcation structure of positive solutions of (3.1) (equivalently (3.2)), together with their asymptotic behavior as the nonlinear diffusion parameter $\alpha\to\infty$.
@@
The curve $\lambda=\sigma_{1}(b\mu\mathrm{1}\hskip-2.5pt\mathrm{l}_{\Omega_{1}},\Omega)$ for $\mu\geq 0$ represents the set of parameter values $(\lambda,\mu)$ where positive solutions bifurcate from the semitrivial solution $(0,\mu)$; this fact will be established in Theorem 3.2 below. The following lemma, derived from [3, Theorem 2.1] (see also [29, Lemma 2.1]), shows that the curve $\lambda=\sigma_{1}(b\mu\mathrm{1}\hskip-2.5pt\mathrm{l}_{\Omega_{1}},\Omega)$ is monotone increasing and uniformly bounded for $\mu\geq 0$, see Figure 2.
@@
In [17, 18], the effects of spatial heterogeneities on the bifurcation branches of coexistence steady states in saturation predator–prey models were investigated, and bifurcation branches with a similar qualitative shape were also observed. However, in that setting, additional bifurcation points appear along those branches, depending on the spatial modes, as shown in [10].
@@ 4.1 A priori estimates and the nonexistence region for positive stationary solutions
Let $x_{0}\in\overline{\Omega}_{1}$ be a maximum point of $v$, so that $v(x_{0})=\max_{\overline{\Omega}_{1}}v$. Then applying the maximum principle (e.g., [25, Proposition 2.1]) to the second equation of (3.2), we have
@@
In addition, we note that $\sigma_{1}(\xi\mathrm{1}\hskip-2.5pt\mathrm{l}_{\Omega_{1}})$ is continuously differentiable for $\xi>0$. The proof of the $C^{1}$-regularity can be found, for instance, in Yamada [42, Proposition 1.1 (iii)]. In his work, a continuous function, depending continuously on the parameter, is introduced as a potential function. Nevertheless, the proof remains valid in the above setting, where the potential is given by the step function $\xi\mathrm{1}\hskip-2.5pt\mathrm{l}_{\Omega_{1}}$. It follows from the chain rule that $K(\lambda,\mu,\alpha)$ is continuously differentiable. Since
@@
We remark that in the case of $\alpha=0$, (3.2) has no positive solution if $\lambda\leq\sigma_{1}(b\mu\mathrm{1}\hskip-2.5pt\mathrm{l}_{\Omega_{1}})$ (see [3, Theorem 2.1] and [29, Lemmas 2.1 and 3.3]).
@@ 4.2 Bifurcation structure of positive stationary solutions
and thus $v_{\infty}=\mu$ in $\overline{\Omega}_{1}$. Then $(\lambda_{\infty},u_{\infty},v_{\infty})=(\sigma_{1}(b\mu\mathrm{1}\hskip-2.5pt\mathrm{l}_{\Omega_{1}}),0,\mu)$ must hold by Lemma 4.5. But this contradicts (4.19) and (4.21). Hence the assertion (4.20) holds true. Thanks to Rabinowitz’s global bifurcation theorem [33], together with its refinement to the unilateral/global alternatives for $C^{1}$ Fredholm maps developed in López-Gómez and Mora-Corral [16], López-Gómez [24], and López-Gómez and Sampedro [20] (see also Shi and Wang [34]), $\varGamma_{1}$ satisfies one of the followings:
@@
may exhibit a discontinuous jump of size $bv_{k}(x)$ across the interface $\partial\Omega_{0}\,(\subset\Omega)$. Nevertheless, even in this situation, one can invoke the generalized Harnack inequality due to Lou and Ni [26, Lemma 3.1], whose proof relies on the classical framework of Stampacchia [35], in particular on the Harnack inequality established there for divergence-form operators with merely bounded measurable coefficients and potential functions in $L^{N/2}(\Omega)$.
@@
Declarations
Conflict of interest
The authors declare that they have no competing interests.
Ethical approval
This article does not contain any studies with human participants or animals performed by any of the authors.
Informed consent
Not applicable.
Data availability
Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.